OpenStudy (anonymous):

solve. please show me how you got this.... I am so lost I need step by step help! -15 (3 + x) = -6x -9

OpenStudy (anonymous):

How far did you get?

OpenStudy (anonymous):

ok first you are goign to distribute the -15 into (3+x) and end up with -45-15 x=-6x-9

OpenStudy (anonymous):

Now we want to get all veribles and constants on oppist sides

OpenStudy (anonymous):

and you end up with 9-45=15x-9x

OpenStudy (anonymous):

simplify you get -36=6x and because we are solving for x we will divide -36 by 6 and end up with x=-6

OpenStudy (anonymous):

I started with -15 * 3 and then -15 * x to get rid of ( ) Thank you so much.... I have a few more probs I am stuck on too if your free....

OpenStudy (anonymous):

I had it started right but didn't know what to do next. That is my problem every time.

OpenStudy (anonymous):

Sure ask away

OpenStudy (anonymous):

2 ( 7x + 5) = 3x -11 (2 - x)

OpenStudy (anonymous):

ok remember first you want to distribute the 2 into (7x+5) and the -11 into the (2-x)

OpenStudy (anonymous):

I got to 14x + 10 = 3x -22 - 11x

OpenStudy (anonymous):

Perfect. Now just get all your x terms on one side, and the non-x terms on the other.

OpenStudy (anonymous):

ok move your simplify both sides

OpenStudy (anonymous):

Oh, except when you multipied the -11 to the -1 it should have be a +11x

OpenStudy (anonymous):

err -11 to the -x rather

OpenStudy (anonymous):

No, the 11 wasn't distributed correctly, the right side should be 3x -22 + 11x.

OpenStudy (anonymous):

so its -11 times - x ??

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

You get 14x + 10 = 14x - 22, if you subtract 14x from each side you get 10 = -22, which is false.

OpenStudy (anonymous):

-11(2-x) = -11(2) + (-11)(-x)

OpenStudy (anonymous):

anyone still here?

OpenStudy (anonymous):

I can give it a go

OpenStudy (anonymous):

2 ( 7x + 5) = 3x -11 (2 - x)

OpenStudy (anonymous):

ok the first thing you want to do is to distribute the 2 through (7x+5)

OpenStudy (anonymous):

I think I got it all the way until 14x + 10 = 3x -22 + 11

OpenStudy (anonymous):

You mean 11x. There's an x there too, but everything else is correct

OpenStudy (anonymous):

put an x on your +11

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

now bring the 3x over with the 11x and your number terms to the other side

OpenStudy (anonymous):

then simplify both sides transpose your varibles and constansts and simplify again

OpenStudy (anonymous):

what @ the 14 x on the left?

OpenStudy (anonymous):

when you move things from one side to another, you have to first change the sign of the thing you're moving. the 3x will become negative

OpenStudy (anonymous):

so 14x - 3x + 10 on the left?

OpenStudy (anonymous):

yep. Now you can take the 10 over to the other side in the same way. change the sign, and then you can switch it over to the other side. from there, just add stuff and let me know what you get

OpenStudy (anonymous):

11x=-32 + 11x

OpenStudy (anonymous):

now -11 from both sides?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how can I subtract 11x from 11 I thought you could only do like terms?

OpenStudy (anonymous):

you meant to write 11x, remember? now you can subtract 11x from it

OpenStudy (anonymous):

11x - 11x ??

OpenStudy (anonymous):

Yea. the 11x-11x will cancel, leaving you with a false statement

OpenStudy (anonymous):

Yea it's a false statement, kinda a trick problem

OpenStudy (anonymous):

k how @ -10 - 3/7 x = 5

OpenStudy (anonymous):

Start out the same way getting your varibles and constansts on different sides

OpenStudy (anonymous):

you end up with -3/7x=15

OpenStudy (anonymous):

Have you been taught how to cancel out a fraction?

OpenStudy (anonymous):

I don't know all these rules and when to use them confuse me very BAD!!

OpenStudy (anonymous):

if you move something to one side of the = sign,do you have to move something back over to replace it?

OpenStudy (anonymous):

Inorder to cancel out the fraction we are going to multiply 7*-3 then divide by 7 which will leave us with -3x=15*7, which is -3x=105, now divide both sides by -3 and we end up with x=-35

OpenStudy (anonymous):

No just remember whatever you do to one side you have to do to the other.

OpenStudy (anonymous):

so if we have something like 2+x=1 we are going to subtract both sides by 2 which will give us an answer of x=-1

OpenStudy (anonymous):

k that makes sense

OpenStudy (anonymous):

so does -4/13 x = -20 come out to be.... x = -65 ???

OpenStudy (anonymous):

But like we did above we want to cancel out the fraction so we first need to find a lcm for the equation. This is easy to do because we only have one fraction in the problem. So we first multiply the numerator times the lcm then divide the new numerator by the lcm

OpenStudy (anonymous):

but we have to also multiply the other side by the lcm

OpenStudy (anonymous):

I just realized that can't be right.... it would have to be a positive number wouldn't it?

OpenStudy (anonymous):

correct it would be x=65

OpenStudy (anonymous):

okay.... YAY!!!

OpenStudy (anonymous):

how @ x - 4/11 = 3/5 would it be x = 53/55

OpenStudy (anonymous):

or x = 3/5 ??

OpenStudy (anonymous):

53/55 i think.....

OpenStudy (anonymous):

i agree

OpenStudy (anonymous):

okay you are ahuge help. Thank you...

OpenStudy (anonymous):

No problem best of luck to you

OpenStudy (anonymous):

you know anything @ set builder notation???

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

what would -5x -4 < 3 ( x - 3) be in set builder notation?

OpenStudy (anonymous):

First distribute your 3 like you were before. You want to get some expression for x either greater or less than something else.

OpenStudy (anonymous):

brb, in a few mins

OpenStudy (anonymous):

Ok, back sorry. Did you distribute the 3?